Grade 12Physical Chemistry

density of a 2.05M solution of acetic acid in water is 1.02g/mL.The molality of the solution is???

sajid
9 Years agoGrade 12

4 Answers

Happy Man
9 Years ago
2
Vikas TU
9 Years ago
               Let us find the weight of the acetic acid
   = 2.05 * 60
    =123
The weight of the solution
   =1000* 1.02
    =1020
         Weight of the water =(1020-123)
                                           =897g
               Therefore, Molarity is given as;
                        =2.05 *1000897
                        =2.285molkg-1
Ritik
9 Years ago
We can solve it easily by this formula-d=M[1/m+M`/1000] Where d=density of solution, M=molarity of substance, m= molality of substance, M`=Molecular weight of substance.
Rishi Sharma
6 Years ago
Dear Student,
Please find below the solution to your problem.

Suppose the volume of solution is 1000 ml
1ml=1.2gm
1000ml=1020gm
Number ofr moles =2⋅05M×1lit
=2.05mol
Mass of solute =n×M⋅wt
=2.05×60
=123gm
Solvent mass =1020−123
=897gm
Molality =123​×1000/60×897
​=2.28molkg−1
Thanks and Regards